# On the continuous properties for the 3D incompressible rotating Euler equations

### v1

@@ 2 Preliminaries
$\lesssim\|\partial_{i}u^{j}_{0}\partial_{j}u_{0}^{i}\|_{B^{s-1}_{p,\infty}}$
$\lesssim\|u_{0}\|^{2}_{B^{s}_{p,\infty}}\lesssim 1.$
@@
$b_{n}(x):=2^{-ns}\theta(x)\cos\left(\frac{17}{12}2^{n}x_{1}\right).$
Following similar procedure as that in Theorem 1.3. Then, we have
@@ 6 Discussion and Conclusion
In this paper, we consider the Cauchy problem for the 3D Euler equations with the Coriolis force in the whole space. We first establish the local-in-time existence and uniqueness of solution to this system in $B^{s}_{p,r}(\mathbb{R}^{3})$. Then we prove that the Cauchy problem is ill-posed in two different sense: (1) the solution of this system is not uniformly continuous dependence on the initial data in the same Besov spaces, which extends the recent work of Himonas-Misiołek [19, Comm. Math. Phys., 296, 2010] to the general Besov spaces framework; (2) the solution of this system cannot be Hölder continuous in time variable in the same Besov spaces. In particular, the solution of the system is discontinuous in the weaker Besov spaces at time zero. To the best of our knowledge, our work is the first one addressing the issue on the failure of Hölder continuous in time of solution to the classical Euler equations with(out) the Coriolis force. In this paper, our contributions on the continuous properties for the 3D incompressible rotating Euler equationsare lies in three aspects: by constructing some proper, new and different initial conditions, we showed the non-uniform continuous of the solution map (Theorem 1.2), proved the failure of Hölder regularity of the solution map (Theorem 1.3) and established the discontinuous at $t=0$ of the solution map (Theorem 1.4).
@@
Acknowledgments
J. Li is partially supported by the National Natural Science Foundation of China (No. 12161004), Innovative High end Talent Project in Ganpo Talent Program (No. gpyc20240069), Training Program for Academic and Technical Leaders of Major Disciplines in Ganpo Juncai Support Program (No. 20232BCJ23009). Y. Yu is partially supported by the National Natural Science Foundation of China (No. 12101011). N. Zhu is partially supported by the National Natural Science Foundation of China (No. 12161055), Training Program for Academic and Technical Leaders of Major Disciplines in Ganpo Juncai Support Program (No. 20243BCE51079).
@@ Acknowledgments
Data Availability No data was used for the research described in the article.
Conflict of interest The authors declare that they have no conflict of interest.

### v2

@@ 2 Preliminaries
$\lesssim\|\partial_{i}u^{j}_{0}\partial_{j}u_{0}^{i}\|_{B^{s-1}_{p,\infty}}\lesssim\|u_{0}\|^{2}_{B^{s}_{p,\infty}}\lesssim 1.$
@@
$b_{n}(x):=2^{-n(s+1)}\theta(x)\cos\left(\frac{17}{12}2^{n}x_{1}\right).$
Following similar procedure as that in Theorem 1.3, then we have
@@ 6 Discussion and Conclusion
In this paper, we consider the Cauchy problem for the 3D Euler equations with the Coriolis force in the whole space. We first establish the local-in-time existence and uniqueness of solution to this system in $B^{s}_{p,r}(\mathbb{R}^{3})$. Then we prove that the Cauchy problem is ill-posed in two different sense: (1) the solution of this system is not uniformly continuous dependence on the initial data in the same Besov spaces, which extends the recent work of Himonas-Misiołek [19, Comm. Math. Phys., 296, 2010] to the more general framework of Besov spaces; (2) the solution of this system cannot be Hölder continuous in time variable in the same Besov spaces. In particular, the solution of the system is discontinuous in the weaker Besov spaces at time zero. To the best of our knowledge, our work is the first one addressing the issue on the failure of Hölder continuous in time of solution to the classical Euler equations with(out) the Coriolis force. In this paper, our contribution on the continuous properties for the 3D incompressible rotating Euler equations, unfold in three aspects: by constructing some proper, new and different initial conditions, we showed the non-uniform continuous of the solution map (Theorem 1.2), proved the failure of Hölder regularity of the solution map (Theorem 1.3) and established the discontinuous at $t=0$ of the solution map (Theorem 1.4).
@@
Availability of data and materials
No data was used for the research described in the article.
Conflict of interest
The authors declare that they have no conflict of interest.
Funding
The authors would like to thank the anonymous referees for valuable comments and suggestions which greatly improved the presentation of this paper. J. Li is supported by the National Natural Science Foundation of China (No. 12161004), Innovative High end Talent Project in Ganpo Talent Program (No. gpyc20240069), Training Program for Academic and Technical Leaders of Major Disciplines in Ganpo Juncai Support Program (No. 20232BCJ23009), Jiangxi Provincial Natural Science Foundation (No. 20252BAC210004). N. Zhu is partially supported by the National Natural Science Foundation of China (No. 12161055), Training Program for Academic and Technical Leaders of Major Disciplines in Ganpo Juncai Support Program (No. 20243BCE51079).